# VisionReward: Fine-Grained Multi-Dimensional Human Preference Learning for Image and Video Generation

### v2

@@ Abstract
footnotetext: *Equal contributions. {xjz22, h-y22}@mails.tsinghua.edu.cnfootnotetext: Core contributors: Jiazheng, Yu, Jiale, Yuanming, Jiajun, Yuan, Wenbo and Shen.
@@ 1 Introduction
Visual generative models, including text-to-image [9, 33, 35, 34, 2, 31] and text-to-video [17, 15, 42, 54, 4, 50] generation, have recently experienced rapid developments. Through large-scale pretraining, these models can effectively translate textual descriptions into photorealistic images or temporally coherent videos. To further align them with human preferences, reinforcement learning from human feedback (RLHF) [30]—initially introduced in large language models—has recently been adapted to visual generation tasks [49, 12].
A key bottleneck in applying RLHF to visual generation lies in developing effective visual reward models. Recent studies [49, 22, 47] have explored training reward models to predict human visual preferences, enabling automatic evaluation and preference optimization for visual generative models. For evaluation, reward models function as automated metrics that quantitatively measure the alignment between generated outputs and human preference criteria [23]. For optimization, reward models identify reliable directions for improving visual generation models. Essentially, they can provide feedback in reinforcement learning or generate preference pairs, thus reducing dependence on human annotation [3, 10, 7].
Despite recent progress in reward models (RMs) for visual generation, two primary challenges remain: First, lack of interpretability and risk of unexpected bias. Current RMs [49, 22, 47] for visual generation often suffer from limited interpretability. These models inherently involve complex trade-offs among multiple factors, yet their scoring mechanisms lack transparency regarding how such trade-offs are quantified and balanced. This opacity raises concerns about potential unexpected biases. Though multimodal LLMs like Gemini [41] and GPT-4o [1] enhance interpretability through explainable rating rationales, their general-purpose architectures usually underperform specialized black-box models in fine-grained technical assessments [6]. This raises a key dilemma: how to design new preference prediction methods that allow preference prediction to be interpretable while maintaining accuracy.
Second, lack of effective reward models for video generation. The rapid development of text-to-video generative models has intensified the demand for video reward models, yet this crucial component remains underdeveloped to date. Although image reward models can assess individual frame quality, their frame-level nature inherently neglects essential temporal dependencies in video sequences. While VideoScore [12] has pioneered direct video evaluation through learnable metrics, it still suffers from limitations such as insufficient accuracy in preference prediction and optimization in video generation.
@@
To apply it as a reward model for visual generation models, we propose a multi-dimensional consistent strategy during preference optimization. The goal of this strategy is to mitigate unintended and unquantifiable biases. Specifically, a pair of visual samples is used for preference optimization (e.g., DPO [43]) only if one sample is consistently preferred over the other across all nine dimensions above.
Note that existing approaches [52, 25] have attempted to augment human annotations or expand dimensions of human preferences in visual generation. Different from them, VisionReward defines multi-dimensional human preferences with the goal of 1) disentangling distinct factors to decouple human preferences and 2) thus identifying consistently-preferred samples across all dimensions for preference optimization.
@@
  • For visual preference prediction, VisionReward achieves state-of-the-art performance across multiple benchmarks, while maintaining interpretability via hierarchical diagnostic QA and explicit linear weighting. For instance, VisionReward outperforms VideoScore [12] by 17.2% in accuracy on the human preference prediction task.
@@
2 VisionReward
The complete training process of VisionReward and its application methodology during preference optimization are illustrated in  Fig. 2. The training of VisionReward comprises two synergistic stages:
  • Fine-grained Visual Assessment: we develop a hierarchical framework and fine-tune the visual language model to obtain accurate visual assessment capability.
  • Interpretable Preference Learning: Leveraging the fine-tuned visual language model, we convert its results into binary entries and apply linear regression to learn the optimal weight for predicting preferences.
2.1 Fine-grained Visual Assessment
Hierarchical Diagnostic Framework. When evaluating an image or video, human preferences are often a result of the interplay of multiple factors, necessitating a balance among various considerations. To deconstruct human preferences systematically, we develop a hierarchical diagnostic framework across multiple dimensions (e.g., text alignment, fidelity, safety). Each dimension is operationalized through binary questions requiring yes/no responses. There are 5 dimensions and 61 binary questions for text-to-image, while 9 dimensions and 64 binary questions for text-to-video.
Dataset Annotation. To ensure diversity in our annotated data, we sample from various data sources, preparing 48K images and 33K videos for annotation. Through specialized annotation, we obtain an image dataset containing 3 million question-answer pairs and a video dataset with 2 million pairs (Cf. Appendix  Sec. 7 and  Sec. 10 for details).
Visual Instruction Tuning. Specifically, we use CogVLM2 [18] as the base model for image understanding, and CogVLM2-Video [18] as the base model for video understanding. In terms of data, we have obtained millions of annotated binary question-answering pairs. Initially, we performed a balanced sampling on each binary question by addressing the imbalance between positive and negative examples, ensuring a roughly equal number of positive and negative instances associated with each binary question. Then we use balanced instruction tuning dataset consisting of binary questions to fine-tune base VLM.
2.2 Interpretable Preference Learning
Preference Learning. Based on a series of binary responses (yes or no), we use linear weighted summaries to predict human preferences. This is achieved by mapping yes and no to 1 and -1, respectively, and constructing a feature vector $\mathbf{X}=\{x^{i}\}$, where $i=1,\ldots,n$ and $n$ represents the number of binary questions. We use a set of linear weights to obtain a score that accurately reflects human preferences, expressed as follows:
$\mathrm{score}=\mathbf{x}^{T}\mathbf{w}=x_{1}w_{1}+x_{2}w_{2}+\ldots+x_{n}w_{n}.$ (1)
In order to learn these parameters, we collect a dataset of human preferences $\mathcal{D}=\{(\mathbf{X}_{i},\mathbf{X}_{j})\}$, compute the feature difference for each pair, given by $\Delta\mathbf{X}=\mathbf{X}_{i}-\mathbf{X}_{j}$, and the corresponding label is assigned as $y=1$ or $y=0$ depending on the human preference.
We then perform regression using the logistic regression model based on $\Delta\mathbf{X}$. The predicted value (probability) is given by:
$\hat{y}=\sigma(\Delta\mathbf{x}^{T}\mathbf{w})=\frac{1}{1+e^{-\Delta\mathbf{x}^{T}\mathbf{w}}}.$ (2)
The objective function to be minimized during training is defined as:
$\text{Loss}(\mathbf{w})=$ $-\mathbb{E}\left[y\log\left(\sigma(\Delta\mathbf{x}^{T}\mathbf{w})\right)\right.$
$\left.+(1-y)\log\left(1-\sigma(\Delta\mathbf{x}^{T}\mathbf{w})\right)\right].$ (3)
Through minimizing this objective function, we aim to find the optimal weights $\mathbf{w}$ to predict preferences.
Multi-Dimensional Prediction. Since weights of binary questions are get, we can define score of every question by $\text{reward}(q^{i})=w^{i}*x^{i}$, where $w^{i}$ and $x^{i}$ is the corresponding weight and feature result of question $q^{i}$. For every dimension, we can define the score of certain dimension $d_{k}$ as reward $R(d_{k})$, and $R(d_{k})=\sum\text{reward}(q^{i})$ where $q^{i}\in d_{k}$. It’s clearly that the final reward is the sum of $R(d_{k})$ from all dimensions.
After determining the weights for binary-response questions, we formulate the reward function for each question as:
$\text{reward}(q^{i})=w^{i}\cdot x^{i}$ (4)
where $w^{i}$ denotes the weight and $x^{i}$ represents the corresponding feature value for question $q^{i}$. For each evaluation dimension $d_{k}$, we define its dimensional reward $R(d_{k})$ as the summation of rewards from all constituent questions within that dimension:
$R(d_{k})=\sum_{q^{i}\in d_{k}}\text{reward}(q^{i})$ (5)
The total system reward is subsequently computed as the cumulative sum across all dimensions:
$R_{\text{total}}=\sum_{k}R(d_{k})$ (6)
Upon completion of VisionReward training, the system can execute visual diagnosis and linear weighting through a two-phase framework to predict human preferences.
2.3 Multi-Dimensional Preference Optimization
By calculating dimension-specific scores through intra-dimensional weighting, VisionReward facilitates multi-dimensional preference optimization. To empirically validate the model’s capacity, we leverage Direct Preference Optimization (DPO) for Diffusion Models in our experiments, where VisionReward generates multi-dimensional preference pairs to guide the optimization process while maintaining inter-dimensional balance.
@@ 2.3 Multi-Dimensional Preference Optimization
Preliminary. Given a data distribution $q(x_{0})$, Diffusion models [37, 14, 38] contains forward process and reverse process. Forward process $q(x_{1:T}|x_{0})$ gradually add noise to the data $x_{0}$ and reverse process $p_{\theta}(x_{0:T})$ learns transitions to recover data. Training diffusion model can be performed by evidence lower bound [21, 39]:
@@
$L_{\mathrm{DM}}=\mathbb{E}_{\boldsymbol{x}_{0},\boldsymbol{\epsilon}\sim\mathcal{N}(0,\boldsymbol{I}),t}\left[\left\|\boldsymbol{\epsilon}-\boldsymbol{\epsilon}_{\theta}\left(\boldsymbol{x}_{t},t\right)\right\|_{2}^{2}\right].$ (7)
@@
By learning to recover images, diffusion models are able to generate images that are meaningful to humans. However, just learning on recovering data can not learn to meet human’s preference and needs. Some works have explored to use human preference to guide diffusion to move close to humans. Diffusion-DPO [43] introduces to use preference pairs towards direct preference learning. We denote the “winning” and “losing” samples as $x_{0}^{w},x_{0}^{l}$, and the objective is as follows:
$L(\theta)=-\mathbb{E}_{\left(\boldsymbol{x}_{0}^{w},\boldsymbol{x}_{0}^{l}\right)\sim\mathcal{D},t\sim\mathcal{U}(0,T),\boldsymbol{x}_{t}^{w}\sim q\left(\boldsymbol{x}_{t}^{w}\mid\boldsymbol{x}_{0}^{w}\right),\boldsymbol{x}_{t}^{l}\sim q\left(\boldsymbol{x}_{t}^{l}\mid\boldsymbol{x}_{0}^{l}\right)}$
@@
$\left.\left.-\left(\left\|\boldsymbol{\epsilon}^{l}-\boldsymbol{\epsilon}_{\theta}\left(\boldsymbol{x}_{t}^{l},t\right)\right\|_{2}^{2}-\left\|\boldsymbol{\epsilon}^{l}-\boldsymbol{\epsilon}_{\mathrm{ref}}\left(\boldsymbol{x}_{t}^{l},t\right)\right\|_{2}^{2}\right)\right)\right)$ (8)
where $\lambda_{t}$ is the signal-to-noise ratio.
@@
Figure 3: (a) We sample 10,000 human preference pairs from Pick-a-Pic [22] dataset and analyze score deviations across 18 sub-dimensions (represented by the average yes-proportion of checklist questions within each sub-dimension). (b) We compare score deviations for images generated by SDXL [31] before and after Diffusion-DPO fine-tuning [43], using the same 10,000 prompts.
Challenges. Although Diffusion-DPO has designed a direct preference optimization method to fine-tune text-to-image diffusion models, it suffers from being biased and over-optimized on certain dimensions such as aesthetic but gets somewhat worse, for example, less safe and more trouble about human body. We reproduce the procedure of Diffusion-DPO and test the fine-tuned model on 10,000 sampled prompts from training data, using VisionReward to score images.  Fig. 3 illustrates the changes across various dimensions after applying preference learning using the human preference dataset in which factors are coupled with different trade-offs.
MPO: Insight and Solution. We propose a multi-dimensional preference strategy (see Algorithm 1) to ensure comprehensive enhancement across all dimensions. Following Sec. 2.2, we obtain dimensional reward $R(d_{k})$ and total reward $R_{\text{total}}$ for each generated image. For two images $x^{i}$ and $x^{j}$, we define $R^{i}$ as dominating $R^{j}$ if $R(d_{k}^{i})\geq R(d_{k}^{j})$ holds for every dimension $d_{k}$. The key differences between the MPO strategy and standard DPO are:
  • Standard DPO: During DPO optimization, we generate $m$ images per prompt and directly select the pair with the maximal difference in total reward $R_{\text{total}}$.
  • MPO-enhanced DPO: MPO strategy introduces an additional constraint: we only select pairs $x_{0}^{i}$ and $x_{0}^{j}$ if $R^{i}$ dominates $R^{j}$, then proceed with standard DPO.
We analyze the effects of MPO in Sec. 3.3. The MPO strategy can also be applied to other algorithms, which we leave for future exploration.
Algorithm 1 Multi-Dimensional Preference Optimization (MPO)
1:  Dataset: Prompt set $\mathcal{C}$, question set for reward model $\mathcal{Q}=\left\{q_{1},q_{2},...,q_{n}\right\}$
2:  Input: Diffusion model $p_{\theta}$, reward model $r$, the number of generate images per prompt $m$
3:  Initialization: Dominated set for MPO $\mathcal{D}\leftarrow\emptyset$
4:  for $c\in\mathcal{C}$ do
5:     Sample a set of images $\left\{x_{0}^{1},x_{0}^{2},...,x_{0}^{m}\right\}\sim p_{\theta}(x_{0}|c)$
6:     for $i=1$ to $m$ do
7:        Reward set $R_{i}\leftarrow\emptyset$
8:        for $j=1$ to $n$ do
9:           Add reward $r(q_{j}|x_{0}^{i})$ to $R_{i}$
10:     for $i,j\sim\left\{1,2,...,m\right\}$ do
11:        if $R_{i}$ dominate $R_{j}$ then
12:           Add $\left\{c,x_{0}^{i},x_{0}^{j}\right\}$ to $\mathcal{D}$
13:  for $data\sim\mathcal{D}$ do
14:     Update the gradient $p_{\theta}$ from  Eq. 8
Type Image Video
Content People, Objects, Animals, Story, Human Activity,
Architecture, Landscape, Artificial Scene, Others,
Vehicles, Plants, Food, Natural Animal Activity,
Others, Scenes Physical Phenomena
Challenge Unreal, Style, History, Material, Angle and Lens,
Fine-grained Detail, Emotional Expression,
Color, Famous Character, Color/Tone, Surreal,
Normal, Famous Places, World Knowledge,
Writing, Complex Combo, Special Effects, Text,
Positional, Counting, Spatial Relationship,
Camera Movement,
Logical Consistency,
Style, Temporal Speed
Table 1: Content and Challenge Categories of MonetBench.
@@
HPDv2 [47] MonetBench GenAI-Bench [20] MonetBench
@@
ImageReward [49] 74.0 48.8 56.5 48.4 72.1 55.8 58.4
PickScore [22] 79.8 49.8 57.6 52.4 75.4 57.7 61.6
HPSv2 [47] 83.3 48.4 55.6 49.3 73.0 59.3 62.5
MPS [52] 83.5 44.2 50.7 46.9 67.6 55.8 58.9
@@
GPT-4o [1] 77.5 38.9 52.7 41.8 54.3 45.7 48.3
Gemini [41] 60.7 27.4 55.1 46.9 61.7 52.2 56.8
VQAScore [27] 69.7 49.4 56.5 45.2 68.0 56.1 59.5
VideoScore [12] 76.8 45.8 52.5 47.8 71.4 49.1 54.9
@@
Table 2: Preference accuracy on multiple dataset. Bold denotes the best score within the generative models, while underline signifies the best score among all categories. Tau means taking account of ties [8], and diff∗∗ means dropping ties in labels (we drop ties both in labels and responses for GPT-4o and Gemini in diff∗∗ because too many ties are given by them).
Method Image Video
Composition Quality Fidelity Safety&Emotion Stability Dynamic Physics Preservation
LLaVa 59.9 65.7 59.8 64.4 52.5 53.8 50.6 47.5
CogVLM2 [18] 65.8 67.1 53.1 74.7 49.3 57.1 51.2 47.8
GPT-4o [1] 73.1 62.7 61.9 70.1 57.9 69.1 62.4 58.8
Gemini [41] 69.4 59.9 59.7 74.9 58.1 71.1 58.1 59.6
VisionReward (Ours) 78.8 81.1 80.9 83.9 64.8 75.4 68.1 72.0
Table 3: Accuracy of VisionReward and other vision-language models (VLMs) on vision quality questions constructed from our annotation. We test LLaVA-v1.5-7B [28] for image and LLava-Next-Video-34B [24] for video.
Composition Quality Fidelity Safety
97.9 98.2 98.3 99.1
Stability Dynamic Physics Preservation
97.4 99.9 88.2 99.8
Table 4: Consistency of VisionReward in each dimension.

### v3

@@ 1 Introduction
Visual generative models, including text-to-image (ding2021cogview; ramesh2021zero; saharia2022photorealistic; rombach2022high; betker2023improving; podell2023sdxl) and text-to-video (hong2022cogvideo; ho2022imagen; villegas2022phenaki; opensora; chen2024videocrafter2; yang2024cogvideox) generation, have recently experienced rapid developments. Through large-scale pretraining, these models can effectively translate textual descriptions into photorealistic images or temporally coherent videos. To further align them with human preferences, reinforcement learning from human feedback (RLHF) (ouyang2022training)—initially introduced in large language models—has recently been adapted to visual generation tasks (xu2023imagereward; he2024videoscore).
A key bottleneck in applying RLHF to visual generation lies in developing effective visual reward models. Recent studies (xu2023imagereward; kirstain2023pick; wu2023human) have explored training reward models to predict human visual preferences, enabling automatic evaluation and preference optimization for visual generative models. For evaluation, reward models function as automated metrics that quantitatively measure the alignment between generated outputs and human preference criteria (li2023agiqa). For optimization, reward models identify reliable directions for improving visual generation models. Essentially, they can provide feedback in reinforcement learning or generate preference pairs, thus reducing dependence on human annotation (black2023training; fan2023dpok; clark2023directly).
Despite recent progress in reward models (RMs) for visual generation, two primary challenges remain: First, lack of interpretability and risk of unexpected bias. Current RMs for visual generation often suffer from limited interpretability. These models inherently involve complex trade-offs among multiple factors, yet their scoring mechanisms lack transparency regarding how such trade-offs are performed. This opacity raises concerns about potential unexpected biases. Though multimodal LLMs like Gemini (team2024gemini) and GPT-4o (achiam2023gpt) enhance interpretability through explainable rating rationales, their general-purpose architectures usually underperform specialized black-box models in fine-grained assessments (chen2024mj). This raises a key dilemma: how to design preference prediction method to be interpretable while maintaining accuracy.
Second, lack of effective reward models for video generation. The rapid development of text-to-video generative models has intensified the demand for video reward models. Although image reward models can assess individual frame quality, their frame-level nature inherently neglects essential temporal dependencies in video sequences. While VideoScore (he2024videoscore) has pioneered direct video evaluation through learnable metrics, it still suffers from limitations such as insufficient accuracy in preference prediction and optimization in video generation.
@@
To apply it as a reward model for visual generation models, we propose a multi-dimensional consistent strategy during preference optimization. The goal of this strategy is to mitigate unintended and unquantifiable biases. Specifically, a pair of visual samples is used for preference optimization (e.g., DPO (wallace2024diffusion)) only if one sample is consistently preferred over the other across all dimensions.
@@
  • We design fine-grained multi-dimensional preference annotation and build the most fine-grained dataset which contain 81K samples and 5M binary annotation, which enable the training of VisionReward.
  • For visual preference prediction, VisionReward achieves state-of-the-art performance across multiple benchmarks, while maintaining interpretability via hierarchical diagnostic QA and explicit linear weighting. For instance, VisionReward outperforms VideoScore (he2024videoscore) by 17.2% in accuracy on preference prediction.
@@
Dataset Image Video #Samples #Dimensions #Fine-Grained #Annotation
ImageReward (xu2023imagereward) 9K - - 0.1M
Pick-a-Pic (kirstain2023pick) 38K - - 0.6M
HPDv2 (wu2023human) 430K - - 0.8M
RichHF-18K (liang2024rich) 18K 4 - 0.1M
MPS (zhang2024learning) 608K 4 - 2.4M
VideoScore (he2024videoscore) 38K 5 - 0.2M
VisionReward (Ours) 81K 18–20 61–64 5.0M
Table 1: Comparison of dataset of VisionReward and other datasets.
Figure 2: Illustration of fine-grained multi-dimensional design. (Left) For image: 5 dimensions, 18 sub-dimensions, and 61 binary questions. (Right) For video: 9 dimensions, 20 sub-dimensions, and 64 binary questions.
@@
2 Related Work
Reinforcement Learning from Human Feedback (RLHF) (stiennon2020learning; nakano2021webgpt; ouyang2022training) refers to optimizing models with reinforcement learning based on human feedback, which is also explored in image and video generation.
Preference Learning for Visual Generation. There are many works learning from human preferences, which collect human annotation for text-to-image (xu2023imagereward; kirstain2023pick; wu2023human) and text-to-video (he2024videoscore). Note that existing approaches (zhang2024learning; liang2024rich) have attempted to augment human annotations or expand dimensions of human preferences in visual generation. Different from them, VisionReward defines fine-grained multi-dimensional human preferences with the goal of disentangling distinct factors to decouple human preferences, to build more accurate and interpretable RM.
RLHF for Visual Generation. For visual generation tasks, several works have explored RLHF, optimizing from the gradient (xu2023imagereward; wu2024deep) or using a policy-based RL approach (black2023training; fan2023dpok; clark2023directly). All these methods require a reward model (RM) to provide feedback for online learning. Diffusion-DPO (wallace2024diffusion) has proposed to optimize the diffusion model directly using human-labeled preference data. However, most RLHF methods face the issue of biased-optimization. By employing a multi-dimensional method, VisionReward achieves robust RLHF.
@@ 2 Related Work
Preliminary for Diffusion-DPO. Given a data distribution $q(x_{0})$, Diffusion models (sohl2015deep; ho2020denoising; song2020score) contains forward process and reverse process. Forward process $q(x_{1:T}|x_{0})$ gradually add noise to the data $x_{0}$ and reverse process $p_{\theta}(x_{0:T})$ learns transitions to recover data. Training diffusion model can be performed by evidence lower bound (kingma2021variational; song2021maximum):
@@
$L_{\mathrm{DM}}=\mathbb{E}_{\boldsymbol{x}_{0},\boldsymbol{\epsilon}\sim\mathcal{N}(0,\boldsymbol{I}),t}\left[\left\|\boldsymbol{\epsilon}-\boldsymbol{\epsilon}_{\theta}\left(\boldsymbol{x}_{t},t\right)\right\|_{2}^{2}\right],$ (1)
@@
Diffusion-DPO (wallace2024diffusion) introduces direct preference optimization based on preference pairs. We denote the “win” and “lose” samples as $x_{0}^{w},x_{0}^{l}$, and the objective is as follows:
$\mathcal{L}(\theta)=-\mathbb{E}_{\left(\boldsymbol{x}_{0}^{w},\boldsymbol{x}_{0}^{l}\right)\sim\mathcal{D},t\sim\mathcal{U}(0,T),\boldsymbol{x}_{t}^{w}\sim q\left(\boldsymbol{x}_{t}^{w}\mid\boldsymbol{x}_{0}^{w}\right),\boldsymbol{x}_{t}^{l}\sim q\left(\boldsymbol{x}_{t}^{l}\mid\boldsymbol{x}_{0}^{l}\right)}$
@@
$\left.\left.-\left(\left\|\boldsymbol{\epsilon}^{l}-\boldsymbol{\epsilon}_{\theta}\left(\boldsymbol{x}_{t}^{l},t\right)\right\|_{2}^{2}-\left\|\boldsymbol{\epsilon}^{l}-\boldsymbol{\epsilon}_{\mathrm{ref}}\left(\boldsymbol{x}_{t}^{l},t\right)\right\|_{2}^{2}\right)\right)\right).$ (2)
DPO is ordinarily based on overall preference, which may be biased. VisionReward enables Multi-Dimensional Preference Optimization to enhance it.
3 VisionReward
3.1 VisionReward Annotation
Fine-Grained Design. When evaluating an image or video, human preferences are often a result of the interplay of multiple factors (palmer2013visual; ibarra2017image), necessitating a balance among various considerations. To deconstruct human preferences systematically, we develop a fine-grained multi-dimensional framework, as shown in  Table 1 and  Fig. 2. For each sub-dimension, we set options that vary gradually in degree, and decompose these options into a series of binary questions (Cf.  Tables 16, 17, 18 and 19 in Appendix Section 10 for details).
Dataset Preparation. The dataset for annotation is collected from diverse sources. For images, we sample images from multiple popular datasets, including ImageRewardDB (xu2023imagereward), HPDv2 (wu2023human), and Pick-a-Pic (kirstain2023pick), and obtain 48k images after filtering. For videos, we sample prompts from VidProM (wang2024vidprom) and use multiple text-to-video models to generate videos to annotate. To ensure diversity of prompts, we use Rouge-L (lin-2004-rouge) for initial filtering, follow UniFL (zhang2024uniflimprovestablediffusion) to perform a semantic-based filtering, and use ChatGPT (achiam2023gpt) for data cleaning, finally get 10k prompts. Then we use CogVideoX (yang2024cogvideox), VideoCrafter2 (chen2024videocrafter2) and OpenSora (opensora) to generate 30k videos, sample from Panda-70M (chen2024panda) to get 3k real videos, leading to 33k videos for annotation. More details are provided in Appendix Section 7.1.
Annotation Management. To avoid bias of annotators, our annotation management includes professional management and standard document. Cooperating with a specialized company, we strictly conduct annotation training for annotators, select qualified annotators, and perform quality inspection of annotation results. Our annotation document gives clear definitions and provides more than 10 examples for each judgment, to align the standard among annotators. Due to these efforts, the consistency of annotators in the binary results reaches 89.29% (images) and 89.33% (videos).
Annotation Analysis. Through specialized annotation, we obtain an image dataset containing 48k images and 3 million question-answer pairs, while a video dataset with 33k videos and 2 million pairs. More statistical analysis of the annotation results is in Appendix Section 7.2.
3.2 VisionReward Training
The complete training process of VisionReward and its application methodology during preference optimization are illustrated in  Fig. 3.
Fine-grained Visual Assessment. Specifically, we use CogVLM2 (hong2024cogvlm2) as the base model for image understanding, and CogVLM2-Video (hong2024cogvlm2) as the base model for video understanding. In terms of data, we have obtained millions of annotated binary question-answering pairs. Initially, we performed a balanced sampling on each binary question by addressing the imbalance between positive and negative examples, ensuring a roughly equal number of positive and negative instances associated with each binary question. Then we use balanced instruction tuning dataset consisting of binary questions to fine-tune base VLM.
Interpretable Preference Learning. After trained on fine-grained dataset, VisionReward can be adopt to judge any image or video by giving a series of binary response answers (“yes” or “no”) $\{A_{i}\}_{i=1}^{N}$, where $N$ represents the number of binary questions. We define reward of every binary question as $\{x_{i}\}_{i=1}^{N}$:
$x_{i}=\mathds{1}\left[A_{i}=\text{``yes''}\right].$ (3)
We construct a feature vector $X=(x_{1},\ldots,x_{N})$, and use a set of linear weights $W=(w_{1},\ldots,w_{N})$ to obtain the final reward $R$:
$R=\sum_{i=1}^{N}w_{i}\mathds{1}\left[A_{i}=\text{``yes''}\right].$ (4)
In order to learn linear weights $W$, we collect human preferences for pairs of $\{(X_{i},X_{j})\}$. Specifically, we compute the feature difference for each pair, given by $\Delta X=X_{i}-X_{j}$, and the corresponding label is assigned as $y=1$ or $y=0$ depending on the human preference. We then perform logistic regression $y=\Delta XW^{T}$ to learn linear weights $W$:
$\mathcal{L}(W)=$ $-\mathbb{E}\left[y\log\left(\sigma(\Delta XW^{T})\right)\right$.
$\left.+(1-y)\log\left(1-\sigma(\Delta XW^{T})\right)\right].$ (5)
By calculating dimension-specific scores through intra-dimensional weighting, VisionReward facilitates multi-dimensional preference prediction. We note dimensions as $\{\text{dim}_{k}\}_{k=1}^{K}$ where $\text{dim}_{k}$ contains questions belonging to the dimension. Then we define reward for certain dimension as:
$R({\text{dim}_{k}})=\sum_{i\in\text{dim}_{k}}w_{i}\mathds{1}\left[A_{i}=\text{``yes''}\right].$ (6)
3.3 Multi-Dimensional Preference Optimization
To empirically validate the model’s capacity, we leverage Direct Preference Optimization (DPO) (wallace2024diffusion) for Diffusion Models in our experiments, where VisionReward generates multi-dimensional preference pairs to guide the optimization process while maintaining inter-dimensional balance.
@@ 3.3 Multi-Dimensional Preference Optimization
Figure 4: (a) We sample 10,000 human preference pairs from Pick-a-Pic dataset and analyze score deviations across 18 sub-dimensions (represented by the average yes-proportion of checklist questions within each sub-dimension). (b) We show score deviations for images generated by SDXL after Diffusion-DPO, using the same 10,000 prompts.
Challenges. We replicate the Diffusion-DPO training procedure using SDXL (podell2023sdxl) on the Pick-a-Pic (kirstain2023pick) dataset, employing VisionReward for comprehensive data analysis and model evaluation. As demonstrated in  Fig. 4, both the preference data and optimized model exhibit biases across several fine-grained dimensions. These findings not only underscore VisionReward’s capability for fine-grained analysis but also emphasize the necessity for optimization approaches that account for multi-dimensional representation.
MPO: Insight and Solution. Compared to ordinary DPO method which select pairs using overall preference, we propose MPO-enhanced DPO that take account fine-grained multi-dimensional preference. For reward of two samples $R^{i}$ and $R^{j}$, we define $R^{i}$ as dominating $R^{j}$ if $R^{i}(\text{dim}_{k})\geq R^{j}(\text{dim}_{k})$ holds for every dimension $\text{dim}_{k}$. The key differences between the MPO strategy and standard DPO are:
  • Ordinary DPO: During DPO optimization, we directly select the pair based on the total reward $R$.
  • MPO-enhanced DPO: MPO strategy introduces an additional constraint: we only select pairs that $R^{i}$ dominates $R^{j}$, then proceed with standard DPO.
We analyze the effects of MPO in Section 4.3. The MPO strategy can also be applied to other algorithms, which we leave for future exploration.
@@
HPDv2 MonetBench GenAI-Bench MonetBench
@@
ImageReward (xu2023imagereward) 74.0 48.8 56.5 48.4 72.1 55.8 58.4
PickScore (kirstain2023pick) 79.8 49.8 57.6 52.4 75.4 57.7 61.6
HPSv2 (wu2023human) 83.3 48.4 55.6 49.3 73.0 59.3 62.5
MPS (zhang2024learning) 83.5 44.2 50.7 46.9 67.6 55.8 58.9
@@
GPT-4o (achiam2023gpt) 77.5 38.9 52.7 41.8 54.3 45.7 48.3
Gemini (team2024gemini) 60.7 27.4 55.1 46.9 61.7 52.2 56.8
VQAScore (lin2025evaluating) 69.7 49.4 56.5 45.2 68.0 56.1 59.5
VideoScore (he2024videoscore) 76.8 45.8 52.5 47.8 71.4 49.1 54.9
@@
Table 2: Preference accuracy on multiple dataset. Bold denotes the best score within the generative models, while underline signifies the best score among all categories. Tau means taking account of ties (deutsch2023ties), and diff∗∗ means dropping ties in labels (we drop ties both in labels and responses for GPT-4o and Gemini in diff∗∗ because too many ties are given by them).